OpenStudy (anonymous):

solve for x. 2[sqrt x+2]−3=7

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

The first step is add 3 to both sides to go from \[\Large 2\sqrt{x+2} - 3 = 7\] to \[\Large 2\sqrt{x+2} - 3+3 = 7+3\] \[\Large 2\sqrt{x+2} + 0 = 10\] \[\Large 2\sqrt{x+2} = 10\]

jimthompson5910 (jim_thompson5910):

Then what's next?

OpenStudy (anonymous):

23

OpenStudy (anonymous):

first get rid of single numbers, then square the square root and the other side since you are doing the opposite, then subtract the lonely 2

OpenStudy (anonymous):

10-2?

jimthompson5910 (jim_thompson5910):

not quite, you divide both sides by 2 (to undo the multiplication of 2)

jimthompson5910 (jim_thompson5910):

\[\Large \frac{2\sqrt{x+2}}{2} = \frac{10}{2}\] \[\Large \sqrt{x+2} = 5\]

OpenStudy (anonymous):

wouldd that be the end result?